# Does provable absence of barren plateaus imply classical simulability? Or, why we need to rethink variational quantum computing

### v2

@@ A.1 Small dynamical Lie algebras
where $\bm{\theta}=(\theta_{1},\theta_{2},\ldots)$ are the trainable parameters, and $H_{l}$ Hermitian operators taken from some set of generators $\mathcal{G}$. We note that Eq. (15) can also capture an ansatz with fixed, un-parametrized, layers by setting some of the parameters to be constant. It is known that one can measure the potential expressivity of a circuit by analyzing its dynamical Lie algebra [122, 20, 123]
@@
$\mathcal{C}_{\operatorname{poly}\!\mathfrak{g}}\subset{\rm CSIM}_{\mathrm{QE}}\,.$ (19)
@@ A.1.1 $U(1)$-equivariant circuit
Several proofs of absence of barren plateaus for $U(1)$-equivariant have been presented in the literature [20, 48, 49]. Crucially, they all somewhat rely on the fact that the restriction of the dynamical Lie algebra into the subspace with $k$ Hamming weight can be, at most, of dimension $\mathcal{O}(\binom{n}{k}^{2})$, meaning that a necessary condition for the loss function to not exhibit barren plateaus is that the initial state’s Hamming weight is not too large. Note that bitstrings with Hamming weight $n-k$ can trivially be mapped to bitstrings of Hamming weight $k$ by flipping all bits, and, while our following statements equally apply to Hamming weight $n-k$, we will focus on the case of Hamming weight $k\leqslant n/2$. As such, one finds
@@
$\mathcal{C}_{{\rm U(1)equiv}}\subset{\rm CSIM}_{{\rm QE}}\,.$ (22)
@@ A.1.2 $S_{n}$-equivariant
$\mathcal{C}_{{\rm Sn\,equiv}}\subset{\rm CSIM}_{{\rm QE}}\,.$ (25)
@@ A.2 Matchgate circuit and small modules
$\mathcal{C}_{{\rm match}}\subset{\rm CSIM}_{{\rm QE}}\,.$ (28)
@@ A.3 Shallow hardware efficient ansatz
First, we recall that as shown in Ref. [22]
@@
$\mathcal{C}_{{\rm shallowHEA}}\subset{\rm CSIM}_{{\rm QE}}\,.$ (31)
It follows that no problem based on shallow hardware efficient ansätze with a local measurement can be outside of ${\rm CSIM}_{{\rm QE}}$. We note that this result was originally reported in Ref. [28], with similar ideas explored concurrently in the context of quantum process learning in Ref. [91].
@@ A.3.1 Quantum generative modeling
A.4 Small angle initialization
A recurring theme in this and the following sections will be how to classically simulate the loss function values for effective polynomial subspaces for problem classes in ${\rm CSIM}_{\mathrm{QE}}$ without relying on strong symmetries as in Secs. A.1 & A.2 or small entanglement light cones such as in Sec. A.3. Broadly speaking, one can either forward propagate the initial state $\rho$ (or measurements of it) through the circuit, or backward propagate the measurement operator $O$ in order to estimate the loss $\ell_{\bm{\theta}}(\rho,O)$.
@@ A.4 Small angle initialization
On the flip side, one can also backward propagate the target operator through the quantum circuit, which may be particularly efficient in practice if $O$ is relatively simple in some basis (e.g., the Pauli basis) and one has to collect measurements from a non-trivial initial state $\rho$. This can again be done using tensor networks in the form of MPOs or using methods, such as LOWESA [96, 97], which work using the Pauli transfer matrix formalism [95, 99, 151].
@@
A.4.1 Hamiltonian variational ansatz
@@
[…]
Given that $U(\bm{\theta})$ implements a short time evolution, the adjoint action of $U(\bm{\theta})$ on $O$ lives, with high probability for $\bm{\theta}\sim\mathcal{P}$ , in an effective polynomially-sized subspace of operators acting on $\mathcal{O}(1)$ qubits. Thus, one has
@@
Since the operators in the small subspace are all acting on at most $\mathcal{O}(1)$ qubits, we can utilize standard Pauli classical shadows on the initial state [78]. Then, since the parametrized quantum circuit can effectively be expressed as $U(\bm{\theta})\simeq e^{-iK\tau}$, where $K$ is a $k$-local Hamiltonian ($k\in\mathcal{O}(1)$) and $\tau\in\mathcal{O}(1/n)$, we know that the entanglement generated by the circuit is bounded, and we can efficiently compute the loss function via tensor networks. Hence,
$\mathcal{C}_{{\rm HVA}}\subset{\rm CSIM}_{\mathrm{QE}}\,.$ (36)
A.4.2 Gaussian initialization in deep variational quantum circuits
@@
[…]
When the parameters are sampled from a Gaussian distribution whose width decreases with the circuit depth, then the adjoint action of the unitary on the measurements operator lives, with high probability for $\bm{\theta}\sim\mathcal{P}$, in an effective polynomially-sized subspace. This means that
@@
This trade-off between the range of parameter values and circuit depth renders the quantum circuit effectively simulated with tensor networks in $\mathcal{O}(\operatorname{poly}(n))$ time. Again we might need to perform classical shadow tomography on the initial state $\rho$. Putting these results together leads to
$\mathcal{C}_{{\rm GI}}\subset{\rm CSIM}_{\mathrm{QE}}\,.$ (39)
@@
A.5 Quantum convolutional neural networks
@@
A.5.1 Standard QCNN with tracing-out pooling layers
@@
The proof of absence of barren plateaus for QCNNs can be found in Ref. [13]. Therein, the authors showed that if all of the two-qubit gates in the circuit form independent local $2$-designs, then the loss function will not exhibit a barren plateau provided that the reduced states of $\rho$ onto any pair of qubits are not exponentially close to the identity (e.g., if the initial state follows an area law of entanglement). Hence, we can find that
[…]
@@
While this equation indicates that the adjoint action allows us to reach exponentially many Pauli operators, this does not mean that the adjoint action will typically lead to exponentially large subspaces. Indeed, one can verify that, with high probability for $\bm{\theta}\sim\mathcal{P}$, $U^{\dagger}(\bm{\theta})OU(\bm{\theta})$ will live in an effective polynomially large subspace whose basis is composed of Pauli of operators acting on at most $\mathcal{O}(1)$ qubits. Thus, with high probability for some $\bm{\theta}\sim\mathcal{P}$, the QCNNs adjoint action over the measurement leads to an effective polynomially-sized subspace, and so
@@
Given that the small subspace contains operators acting on a small number of qubits, we can again perform standard Pauli classical shadows on the initial state [78] to obtain a classical description. Then, since the circuit is composed of local gates, and contains at most $\mathcal{O}(\log(n))$ layers, we can simulate the loss via tensor networks or any of the techniques outlined at the beginning of Sec. A.4. High-fidelity approximations of $O(\bm{\theta})$ likely require polynomial MPO bond dimension, and algorithms such as LOWESA could employ a simple truncation based on operator weight which has for example been used in Refs. [151, 97]. Thus, we find
$\mathcal{C}_{{\rm QCNN}}\subset{\rm CSIM}_{{\rm QE}}\,.$ (43)
A.5.2 Precision-limited QCNN with controlled measurements pooling layers
Next we consider a class of precision-limited QCNN with controlled measurements in the pooling layers, which we denote as $\mathcal{C}_{{\rm pl-QCNN}}$. Here, the circuit is a QCNN composed of layers of two-qubit gates acting on nearest neighbors which are interleaved with pooling layers where half of the qubits are measured, with the measurement outcomes controlling single-qubit unitaries. We take $\rho$ to be an $n$-qubit state preparable by a circuit with $\mathcal{O}(\operatorname{poly}(n))$ gates when acting on the all-zero state, and $O$ some Pauli operator acting on the last two qubits of the QCNN. All of the two-qubit gates in the circuit are parametrized. We also assume that the QCNN is precision-limited, in the sense that the parameters in the gates can only take a number of discrete values $d$ in $\mathcal{O}(\operatorname{poly}(n))$ (i.e., $\mathcal{P}$ is discrete).
While Ref. [13] did not explicitly prove that this form of precision-limited QCNN avoids barren plateaus, one can intuitively see that it will not concentrate exponentially from the fact that it is less expressive than a non-precision standard QCNN. Since decreasing expressiveness generally makes the QCNN less concentrated [21], we expect that the losses in $\mathcal{C}_{{\rm pl-QCNN}}$ will be in $\overline{{\rm BP}}$ for area law initial states too. Moreover, here we also expect that for some randomly sampled $\bm{\theta}$ from the discrete $\mathcal{P}$ the Heisenberg evolved operator will also be contained with high probability on the polynomially small subspace of Pauli operators acting on at most $\mathcal{O}(\log(n))$ qubits.
Note that while simulation is only available with high probability the losses in $\mathcal{C}_{{\rm QCNN}}$, we can actually efficiently simulate those in $\mathcal{C}_{{\rm pl-QCNN}}$ for all parameter values. This follows from the fact that the local observable $O$ conjugated by a QCNN always has bounded Hilbert–Schmidt norm, and since there are $\mathcal{O}(\exp\left(\operatorname{poly}(n)\right))$ parameter settings of the QCNN at this precision, one can use Clifford shadow measurement [78] to obtain an efficient classical description of $\rho$. Such shadows can then be combined with tensor networks method to prove that
$\mathcal{C}_{{\rm pl-QCNN}}\subset{\rm CSIM}_{\mathrm{QE}}\,.$ (44)
More formally the following theorem holds.
Theorem 1 (Classical simulability of QCNNs at $\operatorname{poly}\left(n\right)$ precision).
Consider a sequence of quantum convolutional neural network $U_{n}$ on $n$ qubits with $\operatorname{O}\left(n^{d}\right)$ parameters implemented at $\operatorname{O}\left(\operatorname{poly}\left(n\right)\right)$ bits of precision. Assume the loss of each QCNN is a function of the measurement of a $k$-local observable $O$ to additive error $\epsilon$ where $k=\operatorname{O}\left(1\right)$. Let $\rho$ be an input to $U_{n}$. Using
@@ A.5.2 Precision-limited QCNN with controlled measurements pooling layers
$N=\mathcal{O}\left(\epsilon^{-2}n^{p}\log\left(\delta^{-1}\right)\right)$ (45)
@@
copies of $\rho$, with probability at least $1-\delta$, one can construct a classical shadow representation of $\rho$ that allows one to estimate the loss to additive error $\epsilon$ over all $M$ parameter settings. This classical shadow can be constructed with no knowledge of $U_{n}$ or $O$.
@@
Fix $\epsilon,\delta=\operatorname{\Theta}\left(1\right)$. Consider a classical shadows description of $\rho$ using the random Clifford basis protocol of Ref. [78], built using
$N=\frac{300\left\lVert O\right\rVert_{\infty}^{2}}{\epsilon^{2}}2^{k}\left(n^{d}+n-k+\log\left(2\delta^{-1}\right)\right)=\operatorname{poly}\left(n\right)$ (46)
@@
$M=2^{n^{d}}2^{n-k}$ (47)
observables with Hilbert–Schmidt norm upper-bounded by $\left\lVert O\right\rVert_{\infty}2^{k/2}$ to additive error $\epsilon$ in classical time $\operatorname{poly}\left(N\right)$. All that remains is demonstrating that there are $M$ observables being measured with Hilbert–Schmidt norm upper-bounded by $\left\lVert O\right\rVert_{\infty}2^{k/2}$ by this QCNN at this precision.
Consider the set of $2^{n-k}$ observables:
$O_{m_{1},\ldots,m_{n-k}}=O\otimes\ket{m_{1}}\bra{m_{1}}\otimes\ldots\otimes\ket{m_{n-k}}\otimes\bra{m_{n-k}}.$ | . (48)
Note that the Hilbert–Schmidt norm of each $O_{m_{1},\ldots,m_{n-k}}$ is upper-bounded by $\left\lVert O\right\rVert_{\infty}2^{k/2}$. There are thus $2^{n-k+n^{d}}$ operators of the form:
$U^{\dagger}O_{m_{1},\ldots,m_{n-k}}U$ (49)
over all parameter settings of $U$ with Hilbert–Schmidt norm upper-bounded by $K$. ∎
We note that one could also plausibly extend these arguments to continuous parametrizations via the covering set arguments employed in Refs. [153, 91].
Appendix B Examples of non-concentrated but also non-simulable loss functions
In this section we showcase an example of a non-concentrated loss function which is not classically simulable. This example showcases that $\overline{{\rm BP}}\neq{\rm CSIM}_{{\rm QE}}$.

### v3

@@ A.1 Small dynamical Lie algebras
where $\bm{\theta}=(\theta_{1},\theta_{2},\ldots)$ are the trainable parameters, and $H_{l}$ Hermitian operators taken from some set of generators $\mathcal{G}$. We note that Eq. (15) can also capture an ansatz with fixed, un-parametrized, layers by setting some of the parameters to be constant. It is known that one can measure the potential expressivity of a circuit by analyzing its dynamical Lie algebra [133, 22, 134, 135, 136, 103]
@@
$\mathcal{C}_{\operatorname{poly}\!\mathfrak{g}}\subset{\rm QESIM}\,.$ (19)
@@ A.1.1 $U(1)$-equivariant circuit
Several proofs of absence of barren plateaus for $U(1)$-equivariant have been presented in the literature [22, 50, 51]. Crucially, they all somewhat rely on the fact that the restriction of the dynamical Lie algebra into the subspace with $k$ Hamming weight can be, at most, of dimension $\mathcal{O}(\binom{n}{k}^{2})$, meaning that a necessary condition for the loss function to not exhibit barren plateaus is that the initial state’s Hamming weight is not too large. Note that bitstrings with Hamming weight nkn-k can trivially be mapped to bitstrings of Hamming weight $k$ by flipping all bits, and, while our following statements equally apply to Hamming weight nkn-k, we will focus on the case of Hamming weight $k\leqslant n/2$. As such, one finds
@@
$\mathcal{C}_{{\rm U(1)equiv}}\subset{\rm QESIM}\,.$ (22)
@@ A.1.2 $S_{n}$-equivariant
$\mathcal{C}_{{\rm Sn\,equiv}}\subset{\rm QESIM}\,.$ (25)
@@ A.2 Matchgate circuit and small modules
$\mathcal{C}_{{\rm match}}\subset{\rm QESIM}\,.$ (28)
@@ A.3 Shallow hardware efficient ansatz
First, we recall that as shown in Ref. [24], and also at the end of this Supplemental Information, one can prove that
@@
$\mathcal{C}_{{\rm shallowHEA}}\subset{\rm QESIM}\,.$ (31)
It follows that no problem based on shallow hardware efficient ansätze with a local measurement can be outside of $\rm QESIM$}. We note that this result was originally reported in Ref. [30], with similar ideas explored concurrently in the context of quantum process learning in Ref. [110].
@@ A.3.1 Quantum generative modeling
A.4 Generic shallow locally circuits
Here, by generic shallow locally circuits we refer to circuits where local two-qubit gates act according to some arbitrary topology, and where, there is a path that connects the input state with the measurement operator that crosses, at most, $\mathcal{O}(\log(n))$ such gates. This architecture encompasses hardware efficient ansatz in one-dimension, quantum convolutional neural networks, but also more general circuits where gates can act according to a generic topology (i.e, 2D and beyond). Importantly, the fact that the gate connectivity is arbitrary means that light-cone arguments such as ones previously presented for the shallow hardware efficient ansatz do not hold.
Let us then define as $\mathcal{C}_{{\rm slocal}}$ the class of all instances where the parameterized quantum circuit is built from local gates acting on pairs of qubits according to some arbitrary topology. We then take $\rho$ to be an $n$-qubit state preparable by a circuit with $\mathcal{O}(\operatorname{poly}(n))$ gates when acting on the all-zero state, and the measurement operator $O$ to be a local Pauli acting on a qubit at the circuit’s output. We assume that the parameters are sampled such that each gate forms an independent $2$-design over $SU(4)$.
The proof of absence of barren plateaus for these architectures can be derived with the results in [106, 107], but for completeness, we also present a proof at the end of this Supplemental Information (Appendix C). As shown therein, if the initial state is pure and satisfies an area law of entanglement, we find
$\mathcal{C}_{{\rm slocal}}\subset\overline{{\rm BP}}\,,\,\,\text{if {\hbox{\rho}} follows an area law of entanglement}.$ (34)
Moreover, the scrambling nature of the circuit effectively suppresses the contribution of high weight Paulis and the dynamics occur in a subspace composed of Paulis with $\mathcal{O}(1)$-bodyness qubits, so that
$\mathcal{C}_{{\rm slocal}}\subset\mathcal{C}_{{\rm polySub}}\,.$ (35)
For a rigorous proof see Ref. [79]. Then, the component of the initial state therein can be obtained via classical shadows, and the simulation follows as a direct application of the algorithm in [79], showing that
$\mathcal{C}_{{\rm slocal}}\subset{\rm QESIM}\,.$ (36)
A.5 Small angle initialization
A recurring theme in this and the following sections will be how to classically simulate the loss function values for effective polynomial subspaces for problem classes in $\rm QESIM$} without relying on strong symmetries as in Secs. A.1 & A.2 or small entanglement light cones such as in Sec. A.3. Broadly speaking, one can either forward propagate the initial state $\rho$ (or measurements of it) through the circuit, or backward propagate the measurement operator $O$ in order to estimate the loss $\ell_{\bm{\theta}}(\rho,O)$.
@@ A.5 Small angle initialization
On the flip side, one can also backward propagate the target operator through the quantum circuit, which may be particularly efficient in practice if $O$ is relatively simple in some basis (e.g., the Pauli basis) and one has to collect measurements from a non-trivial initial state $\rho$. This can again be done using tensor networks in the form of MPOs or using methods, such as Pauli Propagation [116, 117], which work using the Pauli transfer matrix formalism [115, 119, 153].
@@
A.5.1 Hamiltonian variational ansatz
@@
$\mathcal{C}_{{\rm HVA}}\subset\overline{{\rm BP}}\,,\,\,\text{if {\hbox{\rho}} follows an area law of entanglement}.$ (37)
Given that $U(\bm{\theta})$ implements a short time evolution, the adjoint action of $U(\bm{\theta})$ on $O$ lives, with high probability for $\bm{\theta}\sim\mathcal{P}$ , in an effective polynomially-sized subspace of operators acting on $\mathcal{O}(1)$ qubits (see the proof of Theorem 4 in Ref [80] for a detailed proof). Thus, one has
@@
Since the operators in the small subspace are all acting on at most $\mathcal{O}(1)$ qubits, we can utilize standard Pauli classical shadows on the initial state [95]. Then, since the parametrized quantum circuit can effectively be expressed as $U(\bm{\theta})\simeq e^{-iK\tau}$, where $K$ is a $k$-local Hamiltonian ($k\in\mathcal{O}(1)$) and $\tau\in\mathcal{O}(1/n)$, we know that the entanglement generated by the circuit is bounded, and we can efficiently compute the loss function via tensor networks or Pauli Propagation methods (with time and sample complexity guarantees provided by Theorems 2 and 4 in Ref. [80]). Hence,
$\mathcal{C}_{{\rm HVA}}\subset{\rm QESIM}\,.$ (39)
A.5.2 Gaussian initialization in deep variational quantum circuits
@@
$\mathcal{C}_{{\rm GI}}\subset\overline{{\rm BP}}\,,\,\,\text{if {\hbox{\rho}} follows an area law of entanglement}\,.$ (40)
When the parameters are sampled from a Gaussian distribution whose width decreases with the circuit depth, then the adjoint action of the unitary on the measurements operator lives, with high probability for $\bm{\theta}\sim\mathcal{P}$, in a low magic subspace. Concretely, as proven in Ref. [80], it belongs to an effective subspace composed of operators acting on $\mathcal{O}(\text{poly}(n))$ qubits that can be efficiently classically identified using the small-angle Pauli Propagation algorithm. It follows that
@@
This trade-off between the range of parameter values and circuit depth renders the quantum circuit effectively simulated with low magic simulation algorithms [80, 188]. Again we might need to perform Pauli measurements on the initial state $\rho$. Putting these results together leads to
$\mathcal{C}_{{\rm GI}}\subset{\rm QESIM}\,.$ (42)
@@
A.6 Quantum convolutional neural networks
@@
A.6.1 Standard QCNN with tracing-out pooling layers
@@
The proof of absence of barren plateaus for QCNNs can be found in Ref. [15], but can also be derived with the tools presented at the end of this Supplemental Information (Appendix C). Therein, the authors showed that if all of the two-qubit gates in the circuit form independent local $2$-designs, then the loss function will not exhibit a barren plateau provided that the reduced states of $\rho$ onto any pair of qubits are not exponentially close to the identity (e.g., if the initial state follows an area law of entanglement). Hence, we can find that
$\mathcal{C}_{{\rm QCNN}}\subset\overline{{\rm BP}}\,,\,\,\text{if {\hbox{\rho}} follows an area law of entanglement}.$ (43)
@@
While this equation indicates that the adjoint action allows us to reach exponentially many Pauli operators, this does not mean that the adjoint action will typically lead to exponentially large subspaces. Indeed, one can verify that, with high probability for $\bm{\theta}\sim\mathcal{P}$, $U^{\dagger}(\bm{\theta})OU(\bm{\theta})$ will live in an effective polynomially large subspace whose basis is composed of Pauli of operators acting on at most $\mathcal{O}(1)$ qubits [78, 79]. Thus, with high probability for some $\bm{\theta}\sim\mathcal{P}$, the QCNNs adjoint action over the measurement leads to an effective polynomially-sized subspace, and so
@@
Given that the small subspace contains operators acting on a small number of qubits, we can again perform standard Pauli classical shadows on the initial state [95] to obtain a classical description. Then, since the circuit is composed of local gates, and contains at most $\mathcal{O}(\log(n))$ layers, we can simulate the loss via tensor networks or any of the techniques outlined at the beginning of Sec. A.5. High-fidelity approximations of $O(\bm{\theta})$ likely require polynomial MPO bond dimension, and algorithms such as Pauli Propagation could employ a simple truncation based on operator weight which has for example been used in Refs. [153, 117]. Thus, we find
$\mathcal{C}_{{\rm QCNN}}\subset{\rm QESIM}\,.$ (46)
Indeed, we note that the recent work of [78] has explicitly implemented the previous classical simulation method, effectively dequantizing the quantum convolutional neural network architecture.
A.6.2 QCNN with post-selected pooling layers
Next we consider a class of QCNNs where the loss function depends on the outcomes of the measurements in the pooling layers; we denote this class of QCNNs as $\mathcal{C}_{{\rm pl-QCNN}}$. This setting is typically studied in the context of error-correction, where one wants to learn a circuit that optimizes recovery given some noise channel and logical state [105]. Here, the circuit is a QCNN composed of layers of two-qubit gates acting on nearest neighbors, interleaved with pooling layers where half of the qubits are measured and used to control single-qubit unitaries. While simulation is only available with high probability for the case of $\mathcal{C}_{{\rm QCNN}}$, we claim one can simulate those in $\mathcal{C}_{{\rm pl-QCNN}}$ for all parameter values; that is,
$\mathcal{C}_{{\rm pl-QCNN}}\subset{\rm QESIM}\,.$ (47)
We assume there are $p$ parameterized gates, each a $2$-qubit unitary; we denote the unitary implemented by the QCNN as $U_{\bm{\theta}}$, where $\bm{\theta}\in\mathbb{R}^{p}$. In the post-selected setting the loss is some function of the expectation values of observables of the form:
$|O=\sum_{\bm{m}\in\mathcal{M}\subseteq\left\{0,1\right\}^{n-k}}O_{m_{1},\ldots,m_{n-k}}\otimes\ket{\bm{m}}\bra{\bm{m}}$ (48)
for some subset $\mathcal{M}\subseteq\left\{0,1\right\}^{n-k}$ and $k$-local $O_{m_{1},\ldots,m_{n-k}}$ (assumed to have bounded Hilbert–Schmidt norm). A typical example of this setting is when a QCNN is used to learn the decoder of a quantum error correcting code where, given a circuit $V$ encoding a logical $k$-qubit state $\rho_{L}$, one maximizes the probability of correcting $t$ errors after an error channel $\mathcal{N}$ is applied [105]:
[…]
While QCNNs are defined via real parameters—i.e., their parameters are of infinite precision—by a net argument [155, 110] we can assume that each of the $p$ parameterized gates is drawn from a finite set $S_{p,\epsilon}$ of $2$-qubit gates up to incurring a total error $\epsilon$ in diamond distance. More concretely, Theorem 1 of the Supplemental Information of Ref. [155] states that such a net of cardinality:
$S_{p,\epsilon}=\mathcal{O}\left(\operatorname{poly}\left(\frac{p}{\epsilon}\right)\right)$ (50)
exists. We thus assume this setting without loss of generality.
We now show Eq. (47) via the following theorem.
Theorem 1 (Classical simulability of QCNNs).
Fix $\epsilon,\delta=\operatorname{\Theta}\left(1\right)$. Consider a QCNN with $p$ parameterized $2$-qubit gates drawn from $S_{p,\epsilon}$, and consider:
$Tr\left(OU\rho U^{\dagger}\right),$ (51)
where $O$ is as in Eq. (48), $U$ is a unitary implementing the QCNN parameterized by $p$ elements of $S_{p,\epsilon}$, and $\rho$ is an $n$-qubit state. Using
@@ A.6.2 QCNN with post-selected pooling layers
$N=\mathcal{O}\left(\frac{\left\lvert\mathcal{M}\right\rvert^{2}}{\epsilon^{2}}\left(p\log\left(\frac{p}{\epsilon}\right)+\log\left(\left\lvert\mathcal{M}\right\rvert\right)+\log\left(\delta^{-1}\right)\right)\right)$ (52)
@@
copies of $\rho$, with probability at least $1-\delta$, one can construct a classical shadow representation of $\rho$ that allows one to estimate Eq. (51) to additive error $\epsilon$ over all parameter settings. This classical shadow representation can be constructed with no prior knowledge of $U$ or $O$.
@@
Consider a classical shadows description of $\rho$ using the random Clifford basis protocol of Ref. [95], built using
$N=\frac{300\left\lvert\mathcal{M}\right\rvert^{2}\left\lVert O_{\text{max}}\right\rVert_{\text{F}}^{2}}{\epsilon^{2}}\left(p\ln\left(S_{p,\epsilon}\right)+\ln\left(\left\lvert\mathcal{M}\right\rvert\right)+\log\left(2\delta^{-1}\right)\right)$ (53)
copies of $\rho$. Here, $\left\lVert O_{\text{max}}\right\rVert_{\text{F}}$ is the maximal Hilbert–Schmidt norm of the $O_{m_{1},\ldots,m_{n-k}}$, assumed to be $\mathcal{O}\left(1\right)$. By Ref. [95], this classical shadows description of $\rho$ is of size $\operatorname{poly}\left(N\right)$. Theorem S1 combined with Proposition S1 of [95] guarantees that, with probability at least $1-\delta$, this classical shadows description suffices to estimate the expectation values of
$M=S_{p,\epsilon}^{p}\left\lvert\mathcal{M}\right\rvert$ (54)
observables with Hilbert–Schmidt norm upper-bounded by $\left\lVert O_{\text{max}}\right\rVert_{\text{F}}$ each to additive error $\frac{\epsilon}{\left\lvert\mathcal{M}\right\rvert}$ in classical time $\operatorname{poly}\left(N\right)$. All that remains is demonstrating that there are $M$ observables being measured with Hilbert–Schmidt norm upper-bounded by $\left\lVert O_{\text{max}}\right\rVert_{\text{F}}$ by this QCNN.
@@
[…]
The Hilbert–Schmidt norm of each $O_{m_{1},\ldots,m_{n-k}}$ is upper-bounded by $\left\lVert O_{\text{max}}\right\rVert_{\text{F}}=\mathcal{O}\left(1\right)$ by definition. By additionally counting all $S_{p,\epsilon}^{p}$ choices for $U$, there are thus $S_{p,\epsilon}^{p}\left\lvert\mathcal{M}\right\rvert$ operators of the form:
[…]
A.7 Non-unital noisy-circuits
Let us define as $\mathcal{C}_{{\rm non-u}}$ the class of all instances where the circuit is composed of two-qubit gates on a one-dimensional lattice, interleaved by local (single-qubit) noise, with a final layer of single-qubit gates. Furthermore, we assume that the noise is non-unital and contractive [70]. We take $\rho$ to be a generic $n$-qubit state, and the observable $O$ to be local. We assume that all parameters are sampled such that each gate forms an independent $2$-design.
As shown in [70], the effect of the non-unital noise is to effectively make the circuit shallow by “resetting” the qubits after each layer of parametrized gates. Indeed, as proved in Theorem 1 of  [70], one can only consider the last $\mathcal{O}(\log(n))$ layers of the circuit, after which one can use light-cone arguments to prove that
$\mathcal{C}_{{\rm non-u}}\subset\overline{{\rm BP}}\,.$ (57)
Similarly, since the measurement is local, the same light-cone arguments can be used to show that the Heisenberg-evolved measurement operator will act on operators acting on at most $\mathcal{O}(\log(n))$ qubits. That is,
$\mathcal{C}_{{\rm non-u}}\subset\mathcal{C}_{{\rm polySub}}\,.$ (58)
Finally, by either using the classical simulation algorithms of Ref. [70] in the case of 1D circuits, or those in Refs. [79, 114] for arbitrary topologies, one obtains
$\mathcal{C}_{{\rm slocal}}\subset{\rm QESIM}\,.$ (59)
A.8 Parameterized Dynamic Circuit
Parameterised dynamic circuits make crucial use of feedforward operations. A feedforward operation consists of a measurement on a qubit followed by a conditional gate: If the measurement outcome is 0 then $U_{0}=I$ is applied, if instead it is $1$, then $U_{1}$ is applied where
$U_{1}=\begin{pmatrix}\cos\varphi e^{-i\phi}&-i\sin\varphi\\ -i\sin\varphi&\cos\varphi e^{i\phi}\end{pmatrix}\,$. (60)
Let us define $\mathcal{C}_{{\rm dyn}}$ as the class of all instances where the circuit implements a channel $\mathcal{E}(\bm{\theta})=\mathcal{U}_{d}(\bm{\theta}_{d})\circ\ldots\circ\mathcal{U}_{1}(\bm{\theta}_{1})$, where each channel $\mathcal{U}_{j}(\bm{\theta}_{j})$ is a completely positive, trace preserving map on $n$ qubits. The operations in each $\mathcal{U}_{j}$ can be written as a composition of single-qubit and two-qubit operations in an arbitrary circuit topology. It is further assumed that:
  1. 1.
    Every component $\theta_{i}\in\bm{\theta}$ parameterizes only a single operation in $\mathcal{E}$ (i.e., the parameters are independent and uncorrelated).
  2. 2.
    $\mathcal{E}$ is locally scrambling, i.e., is invariant under single-qubit random gates from a unitary $2$-design after every gate.
  3. 3.
    $\mathcal{E}$ has a constant worst-case feedforward distance. That is, we define $f$ be the shortest path from a qubit measurement to a feedforward operation through the backwards light cone of the measurement and take $f\in O(1)$.
  4. 4.
    $\mathcal{E}$ has an average entangling power of $\alpha$ and average swapping power $\beta\geqslant 0$ [189].
Let $O$ be a $k$-local Hamiltonian with $k\in\log(n)$ and let the initial state be arbitrary.
Then, as shown in Ref. [73], similarly to the case of non-unital noise, the effect of the resets is to effectively make the circuit shallow by “resetting” the qubits after each layer of parametrized gates. However, in contrast to non-unital noise, these resets can be implemented in a controlled manner. Indeed, as proven in Theorem 1 of  [73], one finds that
$\mathcal{C}_{{\rm dyn}}\subset\overline{{\rm BP}}\,.$ (61)
Since these circuits are locally scrambling, the arguments of Ref. [190] also apply here, thus providing bounds on the approximation error by considering only the reduced subspace. In particular, the mid circuit measurements at the core of dynamic circuits can be viewed as controlled two qubit rotations followed by qubit resets and thus can be absorbed into the setting considered in Ref. [190]. Thus, we have
$\mathcal{C}_{{\rm dyn}}\subset\mathcal{C}_{{\rm polySub}}\,$, (62)
and
$\mathcal{C}_{{\rm dyn}}\subset{\rm QESIM}\,.$ (63)
Supp. Info. B Examples of non-concentrated but also non-simulable loss functions
In this section we showcase an example of a non-concentrated loss function which is not classically simulable. This example showcases that $\overline{{\rm BP}}\neq{\rm QESIM}$.